OpenStudy (anonymous):

Simplify: -8 3/4 - (-13 3/8) =

OpenStudy (ash2326):

We have \[-8\frac{3}{4}-(-13\frac 38)\] Do you know how to convert a mixed fraction like \(a\frac bc \) to a normal fraction?

OpenStudy (anonymous):

a little bit

OpenStudy (ash2326):

could you try converting this? \[8\frac 34\]

OpenStudy (anonymous):

-35/4

OpenStudy (anonymous):

i already got the answer, thank you

OpenStudy (ash2326):

did you solve it?

OpenStudy (ash2326):

so fast :O

OpenStudy (anonymous):

Yeah, my mom helped me lol

OpenStudy (ash2326):

ha ha moms are the best :)